OpenStudy (anonymous):

What are the zero(s) of the function f(x) = the quantity of 5 x squared minus 25 x, all over x? help please?

OpenStudy (anonymous):

choices are x=5 x=0 and x=5 x=0 x=-5

OpenStudy (sleepyhead314):

set the numerator equal to zero and solve for x

OpenStudy (anonymous):

i don't know how, can you please help me? @sleepyhead314

OpenStudy (sleepyhead314):

5x^2 - 25x = 0 factor out a 5x can you do that? :)

OpenStudy (anonymous):

i feel so stupid /.\ but what do you mean by factor out? it wasnt in the lesson :x @sleepyhead314

OpenStudy (sleepyhead314):

the opposite of distributive property like for example if I had 5x + 5 I would factor out a 5 and get 5(x+1)

OpenStudy (anonymous):

Okay so it would be 5(x+2) or 5(x+1)^2 ;o ? @sleepyhead314

OpenStudy (sleepyhead314):

5x^2 - 25x (5x)x - (5x)5 (5x)(x - 5)

OpenStudy (anonymous):

how'd you go straight from (5x)x - (5x)5 to (5x)(x - 5) im sorry im not understanding

OpenStudy (sleepyhead314):

distributive property a(b+c) = ab + ac I going ab + ac = a(b+c)

OpenStudy (anonymous):

okay i see now. then once you do that what's the next step ? ;o

OpenStudy (sleepyhead314):

set each part of it equal to zero and solve (5x)(x - 5) = 0 so 5x = 0 and x-5 = 0

OpenStudy (anonymous):

does that mean the answer is 0 then? o-o

OpenStudy (sleepyhead314):

one of them I think there are two answers :3

OpenStudy (anonymous):

x=5 and x=0? ;o

OpenStudy (sleepyhead314):

yes! :)

OpenStudy (anonymous):

thank you so much <3!

OpenStudy (sleepyhead314):

glad I could help :)